# Diffusion-based Generative Multicasting with Intent-aware Semantic Decomposition

### v1

@@ III-B Intent-Aware Generative Multicasting
III-B1 Multicasting the Semantic Map over Shared Wireless Resources
@@
III-B2 Orthogonal Transmission of the User Intended Classes
Each user-intended class is similarly compressed by ${\bf b}_{l}=\mathcal{C}({\bf X}_{l};r_{l})$ to get the corresponding datastream ${\bf b}_{l}$, which is then transmitted to users over orthogonal wireless channels, with $\mathcal{CM}(.)$ representing the modulation and coding operator.
@@ III-B2 Orthogonal Transmission of the User Intended Classes
III-C Adaptive Communication/Computation-aware Semantic Latency Minimization with Multi-Stream Synchronization
@@
The reconstruction and synthesis distortion/perception requirements are measured by $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ metrics, respectively. These can be any distortion or perception metrics, e.g. MS-SSIM [55], LPIPS [56], FID [57], etc. According to the rate-distortion-perception theorem [26, 25], $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ are decreasing functions of the compression rates, i.e. $r_{l},r_{0}$. Minimization of the sum users latency with the intent matrix is thereby formulated as
@@
Input : $[{\bf I}]_{kl},\forall k,l$, $\{h_{k}\}_{k=1}^{K}$, $\{\varepsilon_{kl}^{(r)}\}_{k=1}^{K}$, $\{\varepsilon_{k}^{(s)}\}_{k=1}^{K}$, $\{T_{k}^{g}\}_{k=1}^{K}$, $\{B_{l}\}_{l=0}^{L}$, $\Phi_{r}\left(.\right)$, $\Phi_{s}\left(.\right)$, $P_{T}$, $N_{0}$, $\mathrm{Iter}_{\mathrm{max}}$, $\mathrm{Tol}$.
@@
4Initialize Set: $i=0$, Randomize: $\mathbf{z}_{0}$,$\{\alpha_{k}\}$, $\{\beta_{kl}\}$, $\{\mu_{l}\}$, $\delta$, $\sigma$, Calculate: $\mathcal{E}^{(r)}_{l}=\min\left\{\varepsilon_{kl}^{(r)}\big{|}[{\bf I}]_{kl}\neq 0,k\in[K]\right\},~{}\forall l\in[L]$, $\mathcal{E}^{(s)}=\min\left\{\varepsilon_{k}^{(s)},k\in[K]\right\}$.
@@
IV Simulation Results
IV-A Reconstruction/Synthesis Rate Distortion/Perception Curves
@@
IV-B The Simulation Setup
We consider a generative semantic multicasting setup, where a transmitter equipped with a camera multicasts the street scene images to multiple users based on the multi-user intent. The street scene images are taken randomly from the Cityscapes dataset. Cityscapes is a benchmark for semantic understanding of urban street scenes [59], which includes 8 different semantic categories, namely, construction, object, nature, sky, human, vehicle, flat, and void, and a total of 35 semantic object classes. Each category may include several object classes, e.g., the vehicle category covers “Bicycle”, “Motorcycle”, “Train”, “Trailer”, “Caravan”, and “Car” object classes. Error-free segmentation of the scene images at transmitter is assumed. For now, we assume each user is interested in a random class of the 35, with no two users interested in the same class, i.e. non-overlapped intents. Later in Subsection IV.H, we investigate effects of overlapped intents as well. The wireless transmission parameters are listed in TABLE I. All results are averaged over 6000 Monte Carlo simulations.
TABLE I: Parameter settings
@@ IV-B The Simulation Setup
Path loss at $d_{0}=1$ m, $\epsilon_{o}$30 dB
@@
Noise power density, $N_{0}$174 dBm/Hz
@@
TABLE II: The average source compression rate $r^{*}$, and the total number of transmitted bits $|{\bf b}|$ for various number of users.
@@
IV-C Comparison with Non-Generative and Full Generative Intent-Unaware Multicasting Benchmarks
We compare our proposed intent-aware generative semantic multicasting framework with the following two benchmarks. 1) Non-Generative intent-unaware Multicasting (NGM): In this benchmark, the transmitter compresses the original source signal at rate $r_{NGM}=\Phi_{r}^{-1}\left(\min\left\{\varepsilon_{kl}^{(r)}\big{|}k\in[K],l\in[L]\right\}\right)$, and multicasts it to all the users over shared resources; 2) Full Generative intent-unaware Multicasting (FGM): In this benchmark, the transmitter compresses the semantic map at rate $r_{FGM}=\Phi_{s}^{-1}\left(\min\left\{\varepsilon_{k}^{(s)},k\in[K]\right\}\right)$ and multicasts it to all users over shared resources, where each user then carries out SDM to locally synthesize a signal utilizing its received semantic map.
@@ IV-C Comparison with Non-Generative and Full Generative Intent-Unaware Multicasting Benchmarks
In TABLE II, we report the average multicasting semantic compression rate $r^{*}$ (bpp), and the resulting total number of bits transmitted over the wireless channel $|{\bf b}|=\sum_{l=0}^{L}|{\bf b}_{l}|$. As expected, the $r^{*}$ and $|{\bf b}|$ values increase, with increasing number of users $K$. On the other hand, from the rate-distortion/perception curves in Fig. 3, it is straightforward to show that the compression rate for the FGM and NGM benchmarks is $r_{NGM}=\Phi_{r}^{-1}(0.02850)=0.65804$ and $r_{FGM}=\Phi_{s}^{-1}(0.58705)=0.45118$, respectively. Comparing these values with the compression rate achieved by our proposed framework as reported in TABLE II, we see that $r^{*}<r_{NGM}$ for $K\leq 3$, while for $K>3$ we have $r^{*}>r_{NGM}$. This shows that our proposed framework achieves a smaller compression rate for smaller $K\leq 3$ values, thereby transmitting fewer bits over the wireless channel. In comparison with FGM, we have $r^{*}>r_{FGM}$, i.e. our proposed framework increases the total number of bits communicated over the wireless channel as expected, due to additional transmission of the user-intended signals.
In Fig. 4 we compare the per-user latency curves of the proposed scheme with the benchmarks for various number of users $K$, where $(\varepsilon_{kl}^{(r)},\varepsilon_{k}^{(s)})=(0.02850,0.58705)$. Our proposed framework significantly reduces the per-user latency in comparison with NGM while preserving privacy of the non-intended scene information. Referring to TABLE II, despite increasing the total number of bits for $K>3$, our proposed approach reduces the per-user latency in comparison with the conventional NGM even for large $K$, thereby improving scalability. It becomes clear that the amount of multicasted information, i.e. number of multicasted bits, dominates the per-user latency. This is expected due to the fact that among the $K$ users, the user with the worst channel becomes the bottleneck in multicasting, thereby increasing the latency. Despite increasing the number of transmitted bits, our proposed framework reduces the number of multicasted bits, consequently reducing the per-user latency for a fixed total spectrum resources and transmit power budget. According to Fig. 4, our proposed generative multicasting framework reduces the per-user latency by 16.7%, 15.4%, 14.1%, and 12.8% in comparison with NGM for $K=5,10,15,20$, respectively, when the transmission power budget is fixed at $P_{\text{T}}=100$mW. Equivalently, our proposed framework reduces the transmission power required to achieve a target per-user latency of $13$ms, by 55.6%, 50.0%, 44.4%, and 44.4%, for $K=5,10,15,20$, respectively. Finally, in comparison with FGM, our proposed framework only slightly increases the latency, which is due to transmission of the user-intended signals over orthogonal resources.
@@
Figure 6: The achieved average (a) per-user latency, (b) spectral efficiency, and (c) power ratio of the semantic map versus the transmit power budget, for various number of users.
IV-D Effects of the Transmit Power Budget
@@
IV-E Effects of the Number of Users
@@
IV-F Effects of the Generation Latency
@@
TABLE III: The generation time for (mobile) diffusion models on typical desktop/device AI processors.
@@ 38 TFLOPS
In our proposed, generative semantic multicasting scheme, the receivers may either be conventional users equipped with high performance AI processors (e.g. NVIDIA A100 or H100), or mobile users equipped with mobile AI processors (e.g. Apple A18 or M4). On the high performance processors, where we can afford more FLOPs, we use the lightweight diffusion model [32] to maintain the generation quality, while on the mobile processors, we use the recently released MobileDiffusion [60] to enable running the model on-device. In TABLE III, we report the resulting generation latency $T_{k}^{g}$ for each of these cases. It is concluded that, depending on the deployed model and the device processor, the generation latency is in the range $T_{k}^{g}\in[2,10]$ ms. Hence, we set $T_{1}^{g}=\cdots=T_{K}^{g}=T^{g}$ and plot the ratio of the average per-user latency over the generation latency, i.e. $T^{*}/T^{g}$, for various $T^{g}$ values in this range when $K=10$, in Fig. 8. This ratio decreases when the generation latency or transmit power budget increases, demonstrating that the generation latency dominates the total latency with increasing $T_{k}^{g}$ and $P_{T}$.
Figure 8: The average per-user latency over the generation latency $T^{*}/T^{g}$ for various transmit power $P_{T}$ values, $K=10$.
IV-G Effects of the Reconstruction/Synthesis Quality Requirements
In Fig. 9, we study effects of the reconstruction quality requirement $\varepsilon_{kl}^{(r)}$, on the achieved average per-user latency $T^{*}$, the spectral efficiency $\lambda^{*}$, and the optimum power ratio for multicasting of the semantic map $\gamma^{*}$. As the reconstruction quality requirement is relaxed, the per-user latency and the spectral efficiency is reduced, while a larger portion of the total power budget is allocated to multicasting of the semantic map.
@@ IV-G Effects of the Reconstruction/Synthesis Quality Requirements
IV-H Effects of Overlapped Intents
@@
IV-I Visual Quality of the Retrieved Signals
In this subsection, we demonstrate the visual quality of the retrieved signals for various reconstruction/synthesis quality requirements and its effect on the average per-user latency achieved by our proposed framework. To this end, we illustrate the visual quality for two typical street scenes in Fig. 13 and 14, where the transmit power budget is $P_{\text{T}}=100$ mw, the number of users is $K=10$, and the reconstruction/synthesis quality requirements is fixed at $(\varepsilon_{kl}^{(r)},\varepsilon_{k}^{(s)})=(0.02850,0.58705)$ for all users and classes. The figures show the original scene image, a sample image synthesized locally by the semantic diffusion model, and the final retrieved signals that is partially reconstructed and partially synthesized for three users. It should be noted that, as the random seed and consequently the randomized instance of noise input to the diffusion model is different at various users, the image synthesized locally by semantic guidance is different at various users. In both scenes, user 1 is interested in the “Building” class, while users 2 and 3 are interested in the “Car” and “Road” class, respectively. In this case, the visual quality is rather high and the average per-user latency achieved is $T^{*}=12.824$ ms. Next, we slightly relax the reconstruction/synthesis quality requirements fixing it at $(\varepsilon_{kl}^{(r)},\varepsilon_{k}^{(s)})=(0.06781,0.63050)$ for all users and classes, and provide the visual quality for the same two street scenes in Fig. 15 and 16. The visual quality, while still acceptable, does show some degradation in both scenes, i.e. the reconstructed signal portions are slightly pixelated, and the synthesized portions have slightly lost color contrast. Moreover, the average per-user latency is now reduced to $T^{*}=7.5635$ ms. These results demonstrate high visual quality of the proposed generative semantic multicasting framework, while allowing rather simple adaptation to various user requirements under the prevailing channel conditions.

### v2

@@ II-B Intent-Aware Generative Multicasting
II-B1 Multicasting the Semantic Map over Shared Wireless Resources
@@
II-B2 Orthogonal Transmission of the User Intended Classes
Each user-intended class is similarly compressed by ${\bf b}_{l}=\mathcal{C}({\bf X}_{l};r_{l})$ to get the corresponding datastream ${\bf b}_{l}$, which is then transmitted to users over orthogonal wireless channels [^2] Hybrid (non/)orthogonal multiple access techniques, e.g. RSMA, for the (non/)intended subsignal classes can be adopted in future work to further improve the performance of our proposed generative multicasting framework., with $\mathcal{CM}(.)$ representing the modulation and coding operator.
@@ II-B2 Orthogonal Transmission of the User Intended Classes
II-C Adaptive Communication/Computation-aware Semantic Latency Minimization with Multi-Stream Synchronization
@@
The reconstruction and synthesis distortion/perception requirements are measured by $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ metrics, respectively. These can be any distortion or perception metrics, e.g. MS-SSIM [41], LPIPS [42], FID [43], etc. According to the rate-distortion-perception theorem [12], $\Phi_{r}\left(r_{l}\right)$ and $\Phi_{s}\left(r_{0}\right)$ are decreasing functions of the compression rates, i.e. $r_{l},r_{0}$. Minimization of the sum users latency with the intent matrix is thereby formulated as
@@
Input : ${\bf I}]_{kl},\forall k,l$, $\{h_{k}\}_{k=1}^{K}$, $\{\varepsilon_{kl}^{(r)}\}_{k=1}^{K}$, $\{\varepsilon_{k}^{(s)}\}_{k=1}^{K}$, $\{T_{k}^{g}\}_{k=1}^{K}$, $\{B_{l}\}_{l=0}^{L}$, $\Phi_{r}\left(.\right)$, $\Phi_{s}\left(.\right)$, $P_{T}$, $N_{0}$, $\mathrm{Iter}_{\mathrm{max}}$, $\mathrm{Tol}$.
@@
4Initialize Set: $i=0$, Randomize: $\mathbf{z}_{0}$$\{\alpha_{k}\}$, $\{\beta_{kl}\}$, $\{\mu_{l}\}$, $\delta$, $\sigma$, Calculate: $\mathcal{E}^{(r)}_{l}=\min\left\{\varepsilon_{kl}^{(r)}\big{|}[{\bf I}]_{kl}\neq 0,k\in[K]\right\},~\forall l\in[L]$, $\mathcal{E}^{(s)}=\min\left\{\varepsilon_{k}^{(s)},k\in[K]\right\}$.
@@
Figure 3: Separate source-channel coding workflow for generative semantic multicasting.
II-D Separate Source-Channel Coding Workflow
Fig. 3 demonstrates the general flowchart of the proposed generative semantic multicasting framework. The workflow is composed of two phases: 1) Offline Pre-training, and 2) Online Deployment. This two-phase design follows a source-channel coding separation design, where the “Offline Pre-training” phase of the generative model is agnostic to the user channel conditions $\mathbb{E}|h_{k}|^{2}/N_{0}$, user intents $\mathbf{I}]_{kl}$, and their distortion/perception requirements $\{\varepsilon_{kl}^{(r)}\}_{k=1}^{K}$, $\{\varepsilon_{k}^{(s)}\}_{k=1}^{K}$. Through pre-training, the rate-distortion/perception curves are derived, which are then used for adaptation to users’ channel conditions, intents, and distortion/perception requirements, i.e. C/I/R adaptability, during the “Online Deployment” phase. The C/I/R adaptation is achieved by running the proposed Algorithm 1 during the “Online Deployment” phase without requiring end-to-end retraining. Algorithm 1 inputs the multiuser channel conditions $\mathbb{E}|h_{k}|^{2}/N_{0}$, user intents $\mathbf{I}]_{kl}$, and their distortion/perception requirements $\{\varepsilon_{kl}^{(r)}\}_{k=1}^{K}$, $\{\varepsilon_{k}^{(s)}\}_{k=1}^{K}$, i.e. C/I/R feedback from users, along with the rate-distortion/perception curves derived during pre-training to adaptively determine the per-class compression rates $\{r_{l}^{*}\}$ and transmission powers $\{p_{l}^{*}\}$. We should note that the intent feedback overhead is as small as $\left|{\mathbf{I}}\right|=KL$ bits, which is negligible in practice. Moreover, as Algorithm 1 solves a sparsely activated SQP, it has a complexity of $O(I(L+K)^{e})$, where $I$ is the number of outer iterations depending on the stopping Tol, and the scaling exponent $e\leq 2$ in the worst case, but here scales much better due to the sparsity structure. In our experiments, we observed an empirical scaling of $1\leq e\leq 1.5$, which is affordable as this algorithm is run once upon establishment of the multicast communication, and the results stay valid as long as its inputs, i.e. the users C/I/R, remain unchanged.
III Simulation Results
III-A Reconstruction/Synthesis Rate Distortion/Perception Curves
@@
III-B The Simulation Setup
We consider a generative semantic multicasting setup, where a transmitter equipped with a camera multicasts the street scene images to multiple users based on the multi-user intent. The street scene images are taken randomly from the Cityscapes dataset. Cityscapes is a benchmark for semantic understanding of urban street scenes [45], which includes 8 different semantic categories, namely, construction, object, nature, sky, human, vehicle, flat, and void, and a total of 35 semantic object classes. Each category may include several object classes, e.g., the vehicle category covers “Bicycle”, “Motorcycle”, “Train”, “Trailer”, “Caravan”, and “Car” object classes. Error-free segmentation of the scene images at transmitter is assumed. For now, we assume each user is interested in a random class of the 35, with no two users interested in the same class, i.e. non-overlapped intents. Later in Subsection IV.H, we investigate effects of overlapped intents as well. The wireless transmission parameters are listed in TABLE II. All results are averaged over 6000 Monte Carlo simulations.
TABLE II: Parameter settings
@@ III-B The Simulation Setup
Path loss at $d_{0}=1$ m, $\epsilon_{o}$ -30 dB
@@
Noise power density, $N_{0}$ -174 dBm/Hz
@@
TABLE III: The average source compression rate $r^{*}$, and the total number of transmitted bits $|{\bf b}|$ for various number of users.
@@
III-C Comparison with Non-Generative and Full Generative Intent-Unaware Multicasting Benchmarks
We compare our proposed intent-aware generative semantic multicasting framework with the following two benchmarks:
1) Non-Generative intent-unaware Multicasting (NGM): In this benchmark, the transmitter compresses the source signal at rate $r_{NGM}=\Phi_{r}^{-1}\left(\min\left\{\varepsilon_{kl}^{(r)}\big{|}k\in[K],l\in[L]\right\}\right),$ and multicasts it to all the users over shared resources;
2) Full Generative intent-unaware Multicasting (FGM) [37]: In this benchmark, the transmitter compresses the semantic map at rate $r_{FGM}=\Phi_{s}^{-1}\left(\min\left\{\varepsilon_{k}^{(s)},k\in[K]\right\}\right)$ and multicasts it to all users over shared resources, where each user then carries out SDM according to [37], to locally synthesize a signal utilizing its received semantic map.
@@ III-C Comparison with Non-Generative and Full Generative Intent-Unaware Multicasting Benchmarks
In TABLE III, we report the average multicasting semantic compression rate $r^{*}$ (bpp), and the resulting total number of bits transmitted over the wireless channel $|{\bf b}|=\sum_{l=0}^{L}|{\bf b}_{l}|$. As expected, the $r^{*}$ and $|{\bf b}|$ values increase, with increasing number of users $K$. On the other hand, from the rate-distortion/perception curves in Fig. 4, it is straightforward to show that the compression rate for the FGM and NGM benchmarks is $r_{NGM}=\Phi_{r}^{-1}(0.02850)=0.65804$ and $r_{FGM}=\Phi_{s}^{-1}(0.58705)=0.45118$, respectively. Comparing these values with the compression rate achieved by our proposed framework as reported in TABLE III, we see that $r^{*}<r_{NGM}$ for $K\leq 3$, while for $K>3$ we have $r^{*}>r_{NGM}$. This shows that our proposed framework achieves a smaller compression rate for smaller $K\leq 3$ values, thereby transmitting fewer bits over the wireless channel. In comparison with FGM, we have $r^{*}>r_{FGM}$, i.e. our proposed framework increases the total number of bits communicated over the wireless channel as expected, due to additional transmission of the user-intended signals.
In Fig. 5 we compare the per-user latency curves of the proposed scheme with the benchmarks for various number of users $K$, where $(\varepsilon_{kl}^{(r)},\varepsilon_{k}^{(s)})=(0.02850,0.58705)$. Our proposed framework significantly reduces the per-user latency in comparison with NGM. Referring to TABLE III, despite increasing the total number of bits for $K>3$, our proposed approach reduces the per-user latency in comparison with the conventional NGM even for large $K$, thereby improving scalability. It becomes clear that the amount of multicasted information, i.e. number of multicasted bits, dominates the per-user latency. This is expected due to the fact that among the $K$ users, the user with the worst channel becomes the bottleneck in multicasting, thereby increasing the latency. Despite increasing the number of transmitted bits, our proposed framework reduces the number of multicasted bits, consequently reducing the per-user latency for a fixed total spectrum resources and transmit power budget. According to Fig. 5, our proposed generative multicasting framework reduces the per-user latency by 16.7%, 15.4%, 14.1%, and 12.8% in comparison with NGM for $K=5,10,15,20$, respectively, when the transmission power budget is fixed at $P_{\text{T}}=100$mW. Equivalently, our proposed framework reduces the transmission power required to achieve a target per-user latency of 1313ms, by 55.6%, 50.0%, 44.4%, and 44.4%, for $K=5,10,15,20$, respectively. Finally, in comparison with FGM, our proposed framework only slightly increases the latency, which is due to transmission of the user-intended signals over orthogonal resources.
@@
Figure 7: The achieved average (a) per-user latency, (b) spectral efficiency, and (c) power ratio of the semantic map versus the transmit power budget, for various numbers of users.
III-D Effects of the Transmit Power Budget
@@
III-E Effects of the Number of Users
@@
III-F Effects of the Generation Latency
@@
TABLE IV: The generation time for diffusion models on typical desktop/device AI processors.
@@ 38 TFLOPS
Figure 9: Effects of the generation time and the corresponding communication-computation tradeoffs, $K=10$.
In our proposed, generative semantic multicasting scheme, the receivers may either be conventional users equipped with high performance AI processors (e.g. NVIDIA A100 or H100), or mobile users equipped with mobile AI processors (e.g. Apple A18 or M4). On the high performance processors, where we can afford more FLOPs, we use the lightweight diffusion model [18] to maintain the generation quality, while on the mobile processors, we use the recently released MobileDiffusion [46] to enable running the model on-device. In TABLE IV, we report the resulting generation latency $T_{k}^{g}$ for each of these cases. It is concluded that, depending on the deployed model and the device processor, the generation latency is in the range $T_{k}^{g}\in[2,10]$ ms. Hence, we set $T_{1}^{g}=\cdots=T_{K}^{g}=T^{g}$ and plot the ratio of the average per-user latency over the generation latency, i.e. $T^{*}/T^{g}$, for various $T^{g}$ values in this range when $K=10$, in Fig. 9a. This ratio decreases when the generation latency or transmit power budget increases, demonstrating that the generation latency dominates the total latency with increasing $T_{k}^{g}$ and $P_{T}$. Fig. 9b studies the corresponding communication-computation tradeoffs. For example, according to Fig. 9b, a 2 ms increase in the generation latency (6ms >-> 8ms), can be mitigated by increasing the transmit power budget $P_{T}$ (100mw >-> 200mw) to keep the total latency unchanged at $\approx 17$ms.
III-G Effects of the Reconstruction/Synthesis Quality Requirements
In Fig. 10, we study the effects of the reconstruction quality requirement $\varepsilon_{kl}^{(r)}$, on the achieved average per-user latency $T^{*}$, the spectral efficiency $\lambda^{*}$, and the optimum power ratio for multicasting of the semantic map $\gamma^{*}$. As the reconstruction quality requirement is relaxed, the per-user latency and the spectral efficiency is reduced, while a larger portion of the total power budget is allocated to multicasting of the semantic map.
@@ III-G Effects of the Reconstruction/Synthesis Quality Requirements
III-H Effects of Overlapped Intents
@@
III-I Visual Quality of the Retrieved Signals
In this subsection, we demonstrate the visual quality of the retrieved signals for various reconstruction/synthesis quality requirements and their effect on the average per-user latency achieved by our proposed framework. To demonstrate the effectiveness and generalizability of our proposed framework, we provide simulations both on the Cityscapes and COCO-Stuff datasets.
III-I1 Visual Quality on Cityscapes Street Scenes Dataset
To this end, we illustrate the visual quality for two typical street scenes in Fig. 14 and 15, where the transmit power budget is $P_{\text{T}}=100$ mw, the number of users is $K=10$, and the reconstruction quality requirements is fixed at $(\varepsilon_{kl}^{(r)},\varepsilon_{k}^{(s)})=(0.02850,0.58705)$ for all users and classes. The figures show the original scene image, a sample image synthesized locally by the semantic diffusion model, and the final retrieved signals that is partially reconstructed and partially synthesized for three users. It should be noted that, as the random seed and consequently the randomized instance of noise input to the diffusion model is different at various users, the image synthesized locally by semantic guidance is different at various users. In both scenes, user 1 is interested in the “Building” class, while users 2 and 3 are interested in the “Car” and “Road” class, respectively. In this case, the visual quality is rather high and the average per-user latency achieved is $T^{*}=12.824$ ms. Next, we slightly relax the reconstruction/synthesis quality requirements fixing it at $(\varepsilon_{kl}^{(r)},\varepsilon_{k}^{(s)})=(0.06781,0.63050)$ for all users and classes, and provide the visual quality for the same two street scenes in Fig. 16 and 17. The visual quality, while still acceptable, does show some degradation in both scenes, i.e., the reconstructed signal portions are slightly pixelated, and the synthesized portions have slightly lost color contrast. Moreover, the average per-user latency is now reduced to $T^{*}=7.5635$ ms. These results demonstrate high visual quality of the proposed generative semantic multicasting framework, while allowing rather simple adaptation to various user requirements under the prevailing channel conditions.